OpenStudy (anonymous):

expand the logarithm. state the properties of logarithms used. log z^2 /5

OpenStudy (saifoo.khan):

\[\Large \log \frac{z^2}{5}?\]

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

or is it like this? \[\Large \frac{\log z^2}{5}?\]

OpenStudy (anonymous):

no, like the 1st one

OpenStudy (saifoo.khan):

Okay.

OpenStudy (saifoo.khan):

\[\Large \log (\frac{z^2}{5}) \to \to {\log z^2}-{\log 5} \to \to 2 \log z - \log 5\]

OpenStudy (anonymous):

is that it?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

thanks!!!